INAF LBC+MODS queue observing, 2012 January 19-20

Observer: K. Boutsia, R. Speziali
Telescope Operator: S. Allanson
Support Astronomer: O. Kuhn
Instrument: MODS

Summary

Clear skies at sunset.
We had tecnical problems with the AO secondary ripping (30min lost because we had to repeat an aquisition) and MODS computer hunging (30min lost while Rick was solving the problem). ~1hr lost in total.
We concluded macsj0717 and we started macsj0647 and cosmos-mask1.
The seeing was variable between 1.0''-1.2'' for most of the night, with peacks of 2.0'' during the last hours
Calibration files for all MODS observations can be found in /Repository/20120116 - /Repository/20120117.

Data Summary

Project Objects Filter ExpTime Notes
Nonino macsj0647 dual grating 3.5hr flats+arcs+STD
Nonino macsj0717 dual grating 2.5hr flats+arcs+STD - considered complete
Fiore cosmos dual grating ~2.0hr flats+arcs+STD, the last half hour is diveded in 2 images and the elevation was really low...
STD g191b2 dual grating    
STD BD+75d325 dual grating    
STD Feige34 dual grating    

Details

11:48 Starting up MODS
  • mods1 start gui
  • execMODS ~MODSeng/modsScripts/Support/modsWake.pro
  • modsDisp
  • execMODS bias8K.cal mods1b 0004-0013 / mods1r 0003-0012

01:10 the LDG rotator faulted

01:30 Slewing for pointing and collimation

01:34 acqMODS g191b2b.acq mods1r 00013-0015
  • offsetxy 1.376 8.721 rel
  • seeing in the guider 0.9''

01:42 execMODS g191b2b.obs mods1b 0014 / mods1r 0016
  • the PI here is Pogge ....

01:50 acqMODS macsj0647_mods.acq mods1r 0017-0019
Computed Mask Alignment Offset:
   dX:  -0.097 arcsec
   dY:  -1.115 arcsec
  dPA: -0.1143 degrees

MODS Offset Command:
   offsetpointing -0.1143 -0.097 -1.115 detxy rel

02:08 execMODS macsj0647_mods.obs mods1b 0015-0017 / mods1r 0020-0022
  • we start with 10 aquisitions ...

3:01 GCS interface went down + secondary shell ripped

3:02 Acquisition of the field again acqMODS macsj0647_mods.acq mods1r 0023-0025
  • offsetpointing -0.1091 0.202 -1.607 detxy rel

3:19 execMODS macsj0647_mods.obs mods1b 0018-0028 / mods1r 0026-0036
  • the seeing is getting a bit worse as the wind is picking up (04:42)
  • dimm seeing 1.4'' - seeing on GCS 1.3'' (04:42)
  • gcs seeing still ~1.2''-1.3'' (05:08)
  • we can see many continua (~10) in the slits and in a couple of them there are also emission lines (in the 900sec single exposure)
  • still ~1.2'' of seeing on GCS (05:40)
  • GCS seeing ~1.1'' (06:17)

06:32 acqMODS standard_star.acq mods1r 0037-0039
  • BD+75d325

06:45 execMODS standard_star.obs mods1b 0029 / mods1r 0040
  • counts look good for both channels

06:51 acqMODS macsj0717_lbt_mods_27_IX_2011.acq - ABORT
  • there was a typo in the slit name
  • the coordinates for the guide star were missing!!

07:07 abort to unwrap the telescope!!!

07:09 acqMODS macsj0717_lbt_mods_27_IX_2011.acq mods1r 0041-0043
  • offsetpointing -0.1867 0.597 -1.547 detxy rel

07:20 execMODS macsj0717_lbt_mods_27_IX_2011.obs mods1b 0030-0039 / mods1r 0044-0053
  • in this position we are facing the wind
  • GCS seeing 1.0''-1.1''
  • GCS seeing ~1.0'' (08:47)
  • during this block seeing was varible between 1.0'' and 1.2''.

10:09 acqMODS mods.56.FIORE1.acq mods1r 0054-0056
  • it took 5min to slew ...
  • offsetpointing -0.1837 0.202 -1.028 detxy rel

10:28 execMODS mods.56.FIORE1.obs mods1b 0040-0041 / mods1r 0057-0058
  • we can see the objects through the slits but some of them seem slightly shifted. The problem is that all the ref stars are offset in directions that don't suggest a global offset for the whole mask so that we can twik it and the guiding is a bit jumpy. Check proper motion of ref stars??
  • we go on with the MOS

11:35 the readouts are stuck while it was saving a 30min integration - call Rick
  • the images that it was saving while it hung were mods1b.0040 and mods1r.0057

11:48 We abort the script after Rick's suggestion
  • in theory the images that it was saving are saved on a transfer data disk.
  • Rick managed to recover the 2 images and reset the MODS gui
  • in the meantime we are still guiding so we can send the *obs script without acquiring again (...hopefully)
  • 30min lost but fortunately no images!!

12:00 execMODS mods.56.FIORE1.obs mods1b 0042-0043 / mods1r 0059-0060

12:50 Guiding is not reliable. OB stopped.
  • the images mods1b.0043 and mods1r.0060 were 1000sec long.

12:59 OB restarted but we are entering twilight so we don't know how long we can integrate mods1b 0044 / mods1r 0061
  • we will try to do the 800sec to finish the last integration
  • these last images have 800sec integration

13:16 acqMODS Feige34_std.acq mods1r 0062-0064

13:26 execMODS Feige34_std.obs mods1b 0045 / mods1r 0065
  • it might not be well centered!

13:35 take some MODS exposures testing the centering of the source mods1r 0066

13:37 repeat STD acquisition acqMODS Feige34_std.acq mods1r 0067-0069
  • the GCS is saturated - lost guiding
   Sending offset...
*** ERROR: offsetxy IIF error - OffsetPointing result status: Error, no guide star found in guide box after resumeGuiding
   Offset command returned error - /home/MODSeng/bin/modsAlign aborting

13:46 End of Night

-- KonstantinaBoutsia - 19 Jan 2012
Topic revision: r6 - 21 Jan 2012, KonstantinaBoutsia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback